If the pH of a solution is 5, what is the hydroxide ion concentration?
Answer #1

so [OH]a 1014 109M 10-5 Ansos pH = 5= -log(ihti) do [H]= 10-PH105 m also CH] COH] = 10m - 104 = Kw of water so 10 14 10 3 [